OpenStudy (anonymous):

im stuck ! Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4)

OpenStudy (anonymous):

use the slope-intercept formula: y=mx+b slope (m) should be the same for parallel lines. The variable that changes is b (y-intercept). first write the equation is terms of (-12-4x)/3=y y=(-4/3)x+b plug in the points (-3,4) for x and y then solve for b b=0, therefore the equation for the paralell line is y=(-4/3)x

OpenStudy (anonymous):

but the options are a. y=4/3x+8 b.y= 4/3x+3 c. y=-3/4x+8 d. y=-3/4x+3 the addiction parts is what i dont get.

OpenStudy (anonymous):

ok hold on let me check my math

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got y=-4/3+4

OpenStudy (anonymous):

ah yes the first line simplifies to \[y=\frac{ 4 }{ 3 }x+4\]

OpenStudy (anonymous):

but why would you get rid of the negetive sign ?

OpenStudy (anonymous):

so that means the equation for the parallel line should be \[y=\frac{ 4 }{ 3 }x+b\] plug in (-3,4) gives \[8=b\] \[y=\frac{ 4 }{ 3 }x+8\]

OpenStudy (anonymous):

@thaliluvsya ok hold on let me write up all of the steps for you

OpenStudy (anonymous):

ok thank you because i dont get it.

OpenStudy (anonymous):

\[4x-3y=-12\] \[-3y=-12-4x\] \[y=\frac{ -12-4x }{ -3 }\] \[y=\frac{ -12 }{ -3 }+\frac{ -4x }{ -3 }\] \[y=4+\frac{ 4 }{ 3 }x\]

OpenStudy (anonymous):

this is how I get the line equation for the first line you are given. Do you understand how I get the parallel line?

OpenStudy (anonymous):

i get it just why would you add a negetive sign in front of the 4x behind the 12?

OpenStudy (anonymous):

and would you mind helping me in a few more? just 3